OpenStudy (anonymous):

If \[\lambda x+( 1/x )-1 \] is always non negative for x>0 and \[\lambda > 0\] ,then the least value of \[\lambda \] is ?

OpenStudy (anonymous):

here is my idea start with \[\lambda x+\frac{1}{x}-1>0\] since \(x>0\) can multiply by \(x\) and get \[\lambda x^2+x-1>0\]

OpenStudy (vishweshshrimali5):

-x would be there instead of -1 i think

OpenStudy (anonymous):

parabola that opens up, make sure the vertex is above the \(x\) axis thust a thought though

OpenStudy (vishweshshrimali5):

since you multiplied that with x

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

\(\lambda x^2+x-1\) looks better

OpenStudy (vishweshshrimali5):

\[\lambda x^2+x-x>0\]

OpenStudy (anonymous):

man i still can't write correctly let me try one more time \[\lambda x^2-x+1>0\] how about that one??

OpenStudy (vishweshshrimali5):

\[\lambda x^2>0\]

OpenStudy (anonymous):

@satellite73 Lol... Yes that is correct.

OpenStudy (shubhamsrg):

i'll use y for (lambda) its given yx + 1/x - 1 >= 0 y>= 1/x - 1/x^2 or y>= (x-1)/x^2 we need to find least value of y or in other words max value of (x-1)/x^2 if f(x) = 1/x - 1/x^2 then f'(x) = -1/x^2 + 2/x^3 and f''(x) = 2/x^3 - 6/x^4 for f'(x)= 0, we have x=2 and f''(2) < 0 thus x=2 will be a maxima for (x-1)/x^2 thus min value of y will be 1/4 ..

OpenStudy (vishweshshrimali5):

\[\large{\lambda x+\frac{1}{x}-1>0}\] \[\large{x*(\lambda x+\frac{1}{x}-1)>0}\] \[\large{\lambda x^x+x-x>0}\]

OpenStudy (mathteacher1729):

@shubhamsrg 's calc answer makes good sense. Visually, it checks out too.

OpenStudy (anonymous):

@shubhamsrg How is the least value of y the maximum value of the right side expression? Shouldn't (x-1)/x^2 also be minimum? I get the rest of the answer though.

OpenStudy (anonymous):

@mathteacher1729 do you have any idea why the minimum value of y is the maximum value of (x-1)/x^2 ?

OpenStudy (shubhamsrg):

lets consider the graphs ,,i dont want to figure out what graph RHS will have,,so i take any arbitrary one..we plot x vs lambda  the shaded reigon is where lambda>= RHS so now you must be getting why RHS needs to be max for LHS to be min.. sorry had got DC that time..